### v1

@@ IV Investigating the role of thermal correlations and criticality
In this section, we will examine the scaling factor $\Pi/N$, studying its impact on the system’s behavior for varying values of $N$ (ranging from 10 to 100) and different configurations of $\alpha$, $\beta_{h}$, and $\beta_{c}$. focusing on the behavior near the critical field, including the peaks observed in the paramagnetic and ferromagnetic phases. Additionally, we will discuss how these features depend on $N$ in both the quantum heat engine and refrigerator modes.
@@ IV.1 Critical reaction
We will examine the role of the scaling factor and the behavior at the critical field for $N$ ranging from 10 to 100. We will also analyze the peaks in the paramagnetic phase before the phase transition and the ferromagnetic phase after the transition, as well as their proportionality and dependence on N for the quantum heat engine and refrigerator mode. We intend to use a regression to calculate the exponent $a$ responsible for the growth rate of the scaling factor per spin, proportional to $N^{a}$. We will then examine the behavior of the performance and its trend for small, intermediate, and long chains and the role of thermal gaps and correlations $\alpha$ , to observe the impact of the thermal reservoirs on our system. In the case where $\alpha=0.9$, with $\beta_{h}=0.5$ and $\beta_{c}=1.0$ Fig.14($a$), the obtained exponents are $a=0.4578$ for the critical field, and 0.5154 and 0.46 respectively for the ferromagnetic and paramagnetic phases , with a poor temperature difference between the reservoirs. The scaling factor per particle describes the performance per particle while accounting for the influence of others. We observe that the post-critical and critical exponents are almost identical, with a significant slope for small $N$ (10 to 40) and medium $N$ (40 to 70). This suggests that critical effects slow down the growth and stabilize it in the paramagnetic phase, especially for large $N$ (70 to 100). In the ferromagnetic phase, we see greater growth, with a slight decrease in slope for larger $N$, and the curve is higher than in the critical and paramagnetic cases. For $\alpha=1.2$, with $\beta_{h}=0.2$ and $\beta_{c}=0.9$ Fig.14($b$), the obtained exponents are $a=0.2355$ for the critical field, 0.2308 and 0.2355 for the ferromagnetic and paramagnetic phases , respectively. In this case, $\alpha>1$, which favors more short-range correlations, and the thermal gap is higher compared to the previous case. The exponents are smaller than for $\alpha=0.9$ and are very close to each other, with one of the curves (critical and post-critical) perfectly overlapping The paramagnetic phase exhibits slightly lower growth than the ferromagnetic phase, but both show faster growth with $N$. A similar behavior is observed for small to intermediate $N$, but a notable difference appears for large $N$, where the critical and paramagnetic curves show the onset of pre-saturation the slope increases only marginally between $N=90$ and $N=100$. This indicates that for larger $\alpha$, the system’s performance depends on its size, which significantly amplifies thermal fluctuations and correlations, potentially leading to a more optimal yield. For Fig.14($c$) $\alpha=0.55$, which favors intermediate correlations, with a high thermal gap $\beta_{h}=0.2$ and $\beta_{c}=0.667$ Before the critical field, we observe a curve that is higher than the critical field and the post-phase transition, but it shows a downward behavior indicating underperformance of intermediate chains, which slows down around $N=90$. Near $N=55$ In the ferromagnetic phase, the spins are better aligned, but increasing the number of particles makes the system more sensitive to fluctuations caused by the different types of correlations, which increase the noise and cause the decrease of the scaling factor per spin ($\alpha=-0.1035$). There is an overestimation of the scaling factor per spin of the post-critical field peaks ($\alpha=0.2432$), with an increase of the performance of long chains. At the critical point, there is a slight and slow increase for long chains ($\alpha=0.0504$), indicating a greater difficulty aligning the particles for intermediate correlations before and during the phase transition. In the case of $\alpha=0.2$, with $\beta_{h}=0.6$ and $\beta_{c}=2.0$, Fig.14($d$) the thermal gap is short, low values of $\alpha$ favor long-range correlations, indicating that the fields before the critical point are more efficient. For all chains, performance in the phase transition and after the critical field shows a slight decrease in intermediate N, while the curves for the critical and post-critical peaks demonstrate an increase and outperform of the high chains, which is proportional to $N$.A slight under performance is observed in long chains (from $70$ to $100$) before the phase transition. The scaling factor per spin in the critical fields and after the transition the curves are superimposed for intermediate and shorter chains. This indicates that in long correlation before the critical field there is a superior performance and better spin alignment, resulting in less dissipation and a higher efficiency, there is nonetheless slower growth ($\alpha=0.32826$) compared to postcritical fields ($\alpha=0.9434$) and at the phase transition ($\alpha=0.9413$). We will then move on to the refrigeration mode, which is more prevalent in regions with a small thermal gap. We will examine the influence of the parameter $\alpha$, considering both small and large values, which, respectively, promote long-range and short-range thermal correlations. In the case of $\alpha=0.25$, with $\beta_{h}=1.4$ and $\beta_{c}=1.5$, Fig.14($e$)there is a perfect superposition for intermediate and long chains, with a slight difference for small $N$ for the critical ($\alpha=1.6150$) and postcritical fields ($\alpha=1.2702$). We observe that the performance per particle of the post phase transition for all N is lower than the pre phase transition, which exhibits a significant increase for small chains and saturates at intermediate $N$ ($\alpha=0.5338$), then slightly decreases. In contrast, the critical field reduces the gap for larger chains. It is also noted that the performance before the phase transition is less dependent on the size, which is an effect of long-range correlations. For of $\alpha=1.2$, with $\beta_{h}=0.58$ and $\beta_{c}=0.6$, Fig.14($f$) there is a favoring of short-range correlations. Similarly, there is a perfect superposition between the critical fields and the post-critical peaks ($\alpha=0.4717$) and the the pics before the phase transition increase more slowly ($\alpha=0.4071$), however unlike the previous case they exhibit similar behavior with an increase as the chain size increases, but with a slightly smaller slope for the pre-critical case, which remains higher than the others regardless of $N$. This demonstrates that short-range correlations impose a chain-length dependency.
The scaling factor $N^{\alpha}$ is crucial for understanding how the yield of a quantum thermal engine behaves as the system size increases. The goal is to identify conditions under which increasing $N$ (number of spins or system size) leads to a significant improvement in yield compared to the energy cost. When $\alpha>0$, it means that increasing $N$ results in a superlinear growth of cycle work and engine efficiency, which is a sought after goal to achieve better performance on a large scale. Near a critical point, energy fluctuations and heat capacity can diverge, leading to faster yield growth (higher value of $\alpha$). This is where critical engines can potentially achieve efficiency close to the Carnot point without sacrificing power. For lower values of $\alpha$ (typically far from the critical point), the yield per spin reaches saturation, as spin interactions are not coherent enough to benefit from increasing system size. In conclusion, the results show that the dynamics of thermal efficiency and work per spin depend strongly on the value of $\alpha$, which defines the range of interactions between spins. Temperatures ($\beta_{h}$ and $\beta_{c}$) also influence the nature of transitions, with marked differences in spin thermal sensitivity at different values of $h_{i}$. The observed critical points are generally in agreement with theoretical predictions, although some deviations suggest significant correlation effects in these regimes.
We generally observe better performance when the thermal gaps are very large for $\alpha=1.2$ and $\alpha=0.55$. When short-range correlations are favored, we see a more linear critical behavior that depends more on $N$, particularly with the phase transition. For lower $\alpha$, there is less dependence on the chain length, which can experience a drop of $\pi/N$ for long-range drops ($\alpha=0.55$), or similar behavior before and after the phase transition for large $N$ ($\alpha=0.2$).
@@ V Conclusion
In this study, we investigated the performance of a long-range Ising chain quantum Otto engine and refrigerator under ideal conditions, focusing on the effects of system size $N$, the interaction range parameter $\alpha$, and the temperature ratios $\beta_{h}/\beta_{c}$. Our analysis demonstrates that thermal correlations significantly influence the operational efficiency of these quantum machines, particularly near the critical points of phase transitions. We examined the behavior of all types of correlations for a long chain (N=100). For the quantum heat engine, long-range correlations (small $\alpha$) enhance performance in the precritical region, while short-range correlations (larger $\alpha$) dominate in the post-critical region. Similarly, for the refrigerator, long-range correlations are also favored in the post-critical field, especially in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$.
@@
We then fixed $\alpha$ and varied $N$ to investigate the influence of chain size on performance. For the quantum refrigerator, our results indicate that the cooling efficiency is maximized for smaller system sizes before the phase transition. At the critical point, the behavior of the efficiency $\eta_{R}$ strongly depends on the value of $\alpha$. For long-range correlations, efficiency decreases for short chains but increases for long chains. In contrast, for short-range correlations, $\eta_{R}$ increases, showing better performance for short chains, while long chains exhibit greater stability and reduced dissipation. For the quantum heat engine, we observed relatively stable performance with minimal efficiency loss when the thermal gap is close and transitions occur before the phase transition, where long-range and intermediate correlations show better performance, even as the system size increases. This stability is attributed to the absence of non-adiabatic effects and the slow, quasi-static nature of the transformation. Additionally, the yield $\eta$ differs from that of the refrigerator, as it depends on the chain size for all values of $\alpha$.
Overall, this study highlights the importance of distinguishing between the critical phase and the diamagnetic phase in understanding the performance of quantum heat engines and refrigerators. The ferromagnetic phase, characterized by significant fluctuations and superior performance, contrasts with the diamagnetic phase, which occurs beyond the critical point and exhibits greater stability but inferior performance. These findings underscore the need for a deep understanding of both static and dynamic conditions to effectively optimize the operation of quantum thermal machines.
Appendix A Dynamic of the system
A stationary state is reached in the long-time limit $t\to\infty$. In the following, we will characterize this stationary state through observables. We start with the adjoint version of our master equation, which can be written as:
$\frac{dO_{H}(t)}{dt}=i\{H,O_{H}(t)\}_{-}+\sum_{n,k}\gamma_{n,k}\Big{(}[1-\zeta f_{n}(\omega_{k})]\\ \times[2b_{k}^{\dagger}O_{H}(t)b_{k}-\{b_{k}^{\dagger}b_{k},O_{H}(t)\}_{+}]\\ +f_{n}(\omega_{k})[2b_{k}O_{H}(t)b_{k}^{\dagger}-\{b_{k}b_{k}^{\dagger},O_{H}(t)\}_{+}]\Big{)},$ (9)
where $O_{H}(t)$ denotes the Heisenberg form of a Schrödinger observable $O$.
$\frac{d}{dt}\langle b_{k}^{\dagger}b_{k}\rangle$ $=-2\sum_{n}\gamma_{n,k}\langle b_{k}^{\dagger}b_{k}\rangle+2\sum_{n}\gamma_{n,k}f_{n}(\omega_{k}),$ (64a)
$\frac{d}{dt}\langle b_{k}^{\dagger}b_{q}\rangle$ $=\Big{(}i(\tilde{\omega}_{k}-\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{k}^{\dagger}b_{q}\rangle,$ (64b)
$\frac{d}{dt}\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle$ $=\Big{(}i(\tilde{\omega}_{k}+\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle,$ (64c)
$\frac{d}{dt}\langle b_{q}b_{k}\rangle$ $=\Big{(}-i(\tilde{\omega}_{k}+\tilde{\omega}_{q})-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})\Big{)}\langle b_{q}b_{k}\rangle.$ (64d)
Developing the expressions, we find:
$\langle b_{k}^{\dagger}b_{k}\rangle(t)$ $=\frac{\sum_{n}\gamma_{n,k}f_{n}(\omega_{k})}{\sum_{n}\gamma_{n,k}}\Big{(}1-\exp\big{(}-2\sum_{n}\gamma_{n,k}t\big{)}\Big{)}$
$\quad+\langle b_{k}^{\dagger}b_{k}\rangle_{0}\exp\big{(}-2\sum_{n}\gamma_{n,k}t\big{)},$ (65a)
$\langle b_{k}^{\dagger}b_{q}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}\rangle_{0}\exp\Big{(}i(\tilde{\omega}_{k}-\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65b)
$\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{0}\exp\Big{(}i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65c)
$\langle b_{q}b_{k}\rangle(t)$ $=\langle b_{q}b_{k}\rangle_{0}\exp\Big{(}-i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t-\sum_{n}(\gamma_{n,k}+\gamma_{n,q})t\Big{)},$ (65d)
where the subscript $\langle\cdot\rangle_{0}$ indicates the expectation value at time $t=0$. If $\sum_{n}\gamma_{n,k}\neq 0$, the limit $t\to\infty$ gives:
$\langle b_{k}^{\dagger}b_{q}\rangle_{s}$ $=\delta_{kq}\frac{\sum_{n}\gamma_{n,k}f_{n}(\omega_{k})}{\sum_{n}\gamma_{n,k}},\quad\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{s}=\langle b_{q}b_{k}\rangle_{s}=0.$ (66)
The equations (65) cease to be valid if, for some pair $(k,q)$, $\gamma_{n,k}=\gamma_{n,q}=0$ for all $n$. In that case, we find:
$\langle b_{k}^{\dagger}b_{q}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}\rangle_{0}e^{i(\tilde{\omega}_{k}-\tilde{\omega}_{q})t},$ (67a)
$\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle(t)$ $=\langle b_{k}^{\dagger}b_{q}^{\dagger}\rangle_{0}e^{i(\tilde{\omega}_{k}+\tilde{\omega}_{q})t}.$ (67b)
Appendix B Adiabatic Transformation

### v2

@@ IV Investigating the role of thermal correlations and criticality
In this section, we will examine the scaling factor $\Pi/N$, studying its impact on the system’s behavior for varying values of $N$ (ranging from 10 to 100) and different configurations of $\alpha$, $T_{h}$, and $T_{c}$ . focusing on the behavior near the critical field, including the peaks observed in the paramagnetic and ferromagnetic phases. Additionally, we will discuss how these features depend on $N$ in both the quantum heat engine and refrigerator modes.
@@ IV.1 Critical reaction
We will examine the role of the scaling factor and the behavior at the critical field for $N$ ranging from 10 to 100. We will also analyze the peaks in the paramagnetic phase before the phase transition and the ferromagnetic phase after the transition, as well as their proportionality and dependence on N for the quantum heat engine and refrigerator mode. We intend to use a regression to calculate the exponent $a$ responsible for the growth rate of the scaling factor per spin, proportional to $N^{a}$. We will then examine the behavior of the performance and its trend for small, intermediate, and long chains and the role of thermal gaps and correlations $\alpha$ , to observe the impact of the thermal reservoirs on our system. The goal is to identify conditions under which increasing $N$ (number of spins or system size) leads to a significant improvement in yield compared to the energy cost. When $\alpha>0$, it means that increasing $N$ results in a superlinear growth of cycle work and engine efficiency, to achieve better performance on a large scale.
The parameter $\alpha$ governs the decay of spin interactions, transitioning from a long-range regime ($\alpha\leq 1$) to a short-range regime ($\alpha>1$). At a low $\alpha=0.2$ ($T_{h}=1.67$, $T_{c}=0.5$) , long-range interactions promote strong spin coherence in the ferromagnetic phase ($h_{i}<h_{c}$) for short chains, followed by a sharp decline for larger $N$, resulting in a negative ferromagnetic exponent (a=0.1035a=-0.1035) in Fig. 13-a . This is attributed to a collapse of spin alignment due to heightened susceptibility to thermal and quantum perturbations. Longer chains exhibit modest growth, as long-range interactions enhance fluctuations across a greater number of spins, manifesting more prominently in larger systems with increased resilience post-phase transition. For a high $\alpha=1.2$ ($T_{h}=5.0$, $T_{c}=1.11$) Fig. 13-b , short-range interactions yield small, consistent exponents ($a\approx 0.2355$) across all phases and all $N$, producing stable growth due to the rapid decay of perturbative long-range interactions. This restricts spin coherence, mitigating abrupt losses at $h_{c}$ and favoring the ferromagnetic phase.A pronounced amplification of the critical peak becomes apparent when examining the dependence of performance on system size. As previously discussed, the paramagnetic peak in $\Pi/N$ exhibits a standard linear growth behavior, i.e., $\Pi\sim N$, particularly in the case of long-range interactions characterized by the parameter $\alpha$. In contrast, the peak associated with the critical transition displays a markedly superlinear increase for short-range interactions, while it is suppressed in the presence of long-range, nonlocal interactions.Similarly, the refrigerator exhibits analogous behavior in Fig. 13.c-d , resulting in an increased linearity of $\Pi_{R}/N$ in the ferromagnetic phase for $\alpha=1.2$, with a slope coefficient rising from 0.47 to 0.94(in Fig. 13-d) . This indicates a size-dependent treatment across both phases, with a less pronounced drop compared to the engine mode.In systems with long-range interactions, paramagnetic peaks are enhanced and gradually decrease as the system size increases, eventually saturating to a finite value. This behavior contrasts with that observed for short-range interactions, where the maximum values of the critical peaks follow a clearly superlinear power-law scaling, well described by:
$\Pi_{\text{crit}}^{(R)}/N\sim N^{\alpha},\quad(\alpha>0).$
@@ V Conclusion
In this study, we investigated the performance of a long-range Ising chain quantum Otto engine and refrigerator under ideal conditions, focusing on the effects of system size $N$, the interaction range parameter $\alpha$, and the temperature ratios $T_{c}/T_{h}$. Our analysis demonstrates that thermal correlations significantly influence the operational efficiency of these quantum machines, particularly near the critical points of phase transitions. We examined the behavior of all types of correlations for a long chain (N=100). For the quantum heat engine, long-range correlations (small $\alpha$) enhance performance in the precritical region, while short-range correlations (larger $\alpha$) dominate in the post-critical region. Similarly, for the refrigerator, long-range correlations are also favored in the post-critical field, especially in terms of cooling capacity $Q_{c}/N$ and efficiency $\eta_{R}$.We then fixed $\alpha$ and varied $N$ to investigate the influence of chain size on performance. For the quantum refrigerator, our results indicate that the cooling efficiency is maximized for smaller system sizes before the phase transition. At the critical point, the behavior of the efficiency $\eta_{R}$ strongly depends on the value of $\alpha$. For long-range correlations, efficiency decreases for short chains but increases for long chains. In contrast, for short-range correlations, $\eta_{R}$ increases, showing better performance for short chains, while long chains exhibit greater stability and reduced dissipation. For the quantum heat engine, we observed relatively stable performance with minimal efficiency loss when the thermal gap is close and transitions occur before the phase transition, where long-range and intermediate correlations show better performance, even as the system size increases. This stability is attributed to the absence of non-adiabatic effects and the slow, quasi-static nature of the transformation. Additionally, the yield $\eta$ differs from that of the refrigerator, as it depends on the chain size for all values of $\alpha$.Overall, this study highlights the importance of distinguishing between the critical phase and the paramagnetic phase in understanding the performance of quantum heat engines and refrigerators. The ferromagnetic phase, characterized by significant fluctuations and superior performance, contrasts with the paraagnetic phase, which occurs beyond the critical point and exhibits greater stability but inferior performance. These findings underscore the need for a deep understanding of both static and dynamic conditions to effectively optimize the operation of quantum thermal machines.
Appendix A Long range chain Ising
As a prototypical example of a many-body working substance with, possibly, long-range interactions, we consider a generic model of spinless fermions hopping across the $N$ sites of a linear chain in the presence of pairing interaction, and with a chemical potential $h$. Assuming periodic boundary conditions, the system Hamiltonian reads
$H=$ $-\sum_{j=1}^{N}\sum_{r=1}^{N/2-1}\left[t_{r}c^{\dagger}_{j+r}c_{j}+\Delta_{r}c^{\dagger}_{j+r}c^{\dagger}_{j}+\text{h.c.}\right]$
$-h\sum_{j=1}^{N}\left[1-2c^{\dagger}_{j}c_{j}\right],$ (24)
where $c^{\dagger}_{j}$ and $c_{j}$ are creation and annihilation operators for fermions at site $j$, while $t_{r}$ and $\Delta_{r}$ are the hopping and pairing amplitudes, respectively. We choose their dependence on the intersite distance $r$ according to the power laws
$t_{r}=\frac{1}{N_{\alpha}}\frac{J}{r^{\alpha}},\quad\Delta_{r}=\frac{1}{N_{\alpha}}\frac{J}{r^{\alpha}},$ (25)
with the hopping exponent $\alpha_{1}>0$, the pairing exponent $\alpha_{2}>0$, and $N_{\alpha}=\sum_{r=1}^{N/2}r^{-\alpha}$ the Kac scaling factor, which guarantees extensivity of the energy in the case $\alpha_{i}<1$, with $i=1,2$. Hereafter, we set $J=\Delta=1$ as the energy scale and work in units of $\hbar=k_{B}=1$. In particular, in the short-range case with $\alpha\to\infty$, the relation becomes exact through the Jordan–Wigner mapping anad we descibes the ising transversed field model [66].
Due to the translational invariant nature of the couplings, it is useful to write the Hamiltonian in terms of the momentum-space operators
$\tilde{c}_{k}=e^{-i\frac{\pi}{4}}\frac{1}{\sqrt{N}}\sum_{j=1}^{N}e^{ikj}c_{j},$ (26)
where $k=2\pi n/N$, with $n=-N/2+1,\dots,N/2$ (in the following we will drop the $\tilde{}$ on the $c_{k}$ unless it is ambiguous). Then we obtain
$H=\sum_{k}\left[(h-t_{k})\left(c^{\dagger}_{k}c_{k}-c_{-k}c^{\dagger}_{-k}\right)+\Delta_{k}\left(c^{\dagger}_{k}c^{\dagger}_{-k}+c_{-k}c_{k}\right)\right],$ (27)
where $t_{k}$ and $\Delta_{k}$ are the Fourier transforms of the hopping and pairing amplitudes, respectively, which in the thermodynamic limit may be written as
$t_{k}$ $=\frac{\mathrm{Re}\left[\text{Li}_{\alpha_{1}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{1})},$ (28)
$\Delta_{k}$ $=\frac{\mathrm{Im}\left[\text{Li}_{\alpha_{2}}\left(e^{ik}\right)\right]}{\zeta(\alpha_{2})},$ (29)
where $\text{Li}_{\alpha}(z)$ denotes the polylogarithm and $\zeta(\alpha)$ is the Riemann zeta function. We notice that the Hamiltonian in the Fourier space can be decomposed into the sum of single mode Hamiltonians, introducing $\Psi_{k}=(c_{k},c^{\dagger}_{-k})^{T}$
$H=\sum_{k}\Psi^{\dagger}_{k}H_{k}\Psi_{k},$ (30)
$H_{k}=(h-t_{k})\sigma^{z}_{k}+\Delta_{k}\sigma^{x}_{k},$ (31)
[…]
$u_{k}=\cos\frac{\theta_{k}}{2},\quad v_{k}=\sin\frac{\theta_{k}}{2},$ (32)
where $\theta_{k}=\arctan\left[\Delta_{k}/(h-t_{k})\right]$, to obtain
$H=\sum_{k}\omega_{k}(h)\left(\gamma^{\dagger}_{k}\gamma_{k}-\frac{1}{2}\right),$ (33)
with the spectrum
$\omega_{k}(h)=2\sqrt{(h-t_{k})^{2}+\Delta_{k}^{2}}.$ (34)
Appendix B The Nambu formalism
As we have seen, in the ordered case the Hamiltonian can be diagonalized by a Fourier transformation, reducing the problem to a collection of $2\times 2$ “pseudo-spin-1/2” problems, followed by a Bogoliubov transformation, . In the disordered case, we can proceed similarly, but we cannot reduce ourselves to $2\times 2$ problems in a simple way. By using the Nambu formalism, we define a column vector $\bm{\Psi}$ and its Hermitian conjugate row vector $\bm{\Psi}^{\dagger}$, each of length $2L$, by
$\bm{\Psi}=\begin{pmatrix}\hat{c}_{1}\\ \vdots\\ \hat{c}_{L}\\ \hat{c}^{\dagger}_{1}\\ \vdots\\ \hat{c}^{\dagger}_{L}\end{pmatrix}=\begin{pmatrix}\hat{\mathbf{c}}\\ \hat{\mathbf{c}}^{\dagger}\end{pmatrix}$, (35)
[…]
Warning: Notice that the $\bm{\Psi}$ satisfy quite standard fermionic anti-commutation relations
$\{\Psi_{j},\Psi^{\dagger}_{j^{\prime}}\}=\delta_{j,j^{\prime}},$ (36)
for $j,j^{\prime}=1,\dots,2L$, except that $\{\Psi_{j},\Psi_{j+L}\}=1$ for all $j\leq L$, which brings about certain factors 2 in the Heisenberg’s equations of motion (see later).
It is useful, for later purposes, to introduce the $2L\times 2L$ swap matrix $S$:
$S=\begin{pmatrix}0_{L\times L}&1_{L\times L}\\ 1_{L\times L}&0_{L\times L}\end{pmatrix}$, (37)
in terms of which $\bm{\Psi}^{\dagger}=(S\bm{\Psi})^{T}$.
Consider now a general fermionic quadratic form
$\hat{H}=\sum_{j,j^{\prime}}\left(A_{j^{\prime}j}\hat{c}^{\dagger}_{j^{\prime}}\hat{c}_{j}+A^{*}_{j^{\prime}j}\hat{c}^{\dagger}_{j}\hat{c}_{j^{\prime}}\right)+\sum_{j,j^{\prime}}\left(B_{j^{\prime}j}\hat{c}^{\dagger}_{j^{\prime}}\hat{c}^{\dagger}_{j}+B^{*}_{j^{\prime}j}\hat{c}_{j}\hat{c}_{j^{\prime}}\right),$ (38)
where $A_{j^{\prime}j}=A^{*}_{jj^{\prime}}$, i.e., $A=A^{\dagger}$ is Hermitian, and $B_{jj^{\prime}}=-B_{j^{\prime}j}$, i.e., $B=-B^{T}$ is antisymmetric because $\hat{c}_{j}\hat{c}_{j^{\prime}}$ is antisymmetric under exchange of the two operators, and any symmetric part of $B$ would not contribute. It is simple to verify that $\hat{H}$ can be expressed in terms of $\bm{\Psi}$, omitting an irrelevant constant term $\operatorname{Tr}A$, as:
$\hat{H}=\bm{\Psi}^{\dagger}H\bm{\Psi}=\begin{pmatrix}\hat{\mathbf{c}}^{\dagger},\hat{\mathbf{c}}\end{pmatrix}\begin{pmatrix}A&B\\ -B^{*}&-A^{*}\end{pmatrix}\begin{pmatrix}\hat{\mathbf{c}}\\ \hat{\mathbf{c}}^{\dagger}\end{pmatrix}$. (39)
There is an intrinsic particle-hole symmetry in a fermionic Hamiltonian having this form. This symmetry, further discussed in Sec. 6.1, is connected with the fact that the Hermitian $2L\times 2L$ matrix $H$ satisfies:
$HS=-SH^{*}.$ (40)
In the XY-Ising case, all couplings are real and we have two different fermionic Hamiltonians, one for each parity sector $p=0,1$, which we report here for convenience, using…
Let us consider the eigenvalue problem for a general Hermitian $2L\times 2L$ matrix $H$
$H\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}=\begin{pmatrix}A&B\\ -B^{*}&-A^{*}\end{pmatrix}\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}=\epsilon_{\mu}\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}$, (41)
where $u,v$ are $L$-dimensional column vectors, composing the $2L$-dimensional column vector $\begin{pmatrix}u_{\mu}\\ v_{\mu}\end{pmatrix}$, and the $\mu$ index refers to $\mu$-th eigenvector. By explicitly writing the previous system, we find the so-called Bogoliubov-de Gennes equations:
$\begin{cases}Au_{\mu}+Bv_{\mu}=\epsilon_{\mu}u_{\mu},\\ -B^{*}u_{\mu}-A^{*}v_{\mu}=\epsilon_{\mu}v_{\mu}.\end{cases}$ (42)
It is easy to verify that if $(u_{\mu},v_{\mu})^{T}$ is eigenvector with eigenvalue $\epsilon_{\mu}$, then $(v^{*}_{\mu},u^{*}_{\mu})^{T}$ is an eigenvector with eigenvalue $\epsilon_{\mu}$. We can organize the eigenvectors in a unitary (orthogonal, if the solutions are real) $2L\times 2L$ matrix
$U=\begin{pmatrix}u_{1}&\cdots&u_{L}&v^{*}_{1}&\cdots&v^{*}_{L}\\ v_{1}&\cdots&v_{L}&u^{*}_{1}&\cdots&u^{*}_{L}\end{pmatrix}=\begin{pmatrix}U&V^{*}\\ V&U^{*}\end{pmatrix}$, (43)
$U$ and $V$ being $L\times L$ matrices (real, as we can choose to be, in the Ising case) with the $u_{j}$ and $v_{j}$ as columns. As a consequence:
$U^{\dagger}HU=$ $\begin{pmatrix}\epsilon_{1}&0&\cdots&0&0&0&\cdots&0\\ 0&\epsilon_{2}&\cdots&0&0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\epsilon_{L}&0&0&\cdots&0\\ 0&0&\cdots&0&-\epsilon_{1}&0&\cdots&0\\ 0&0&\cdots&0&0&-\epsilon_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0&0&0&\cdots&-\epsilon_{L}\end{pmatrix}$
$\equiv\text{diag}(\epsilon_{\mu},-\epsilon_{\mu})=E_{\text{diag}}.$ (44)
If we define the new Nambu fermion operators $\bm{\Phi}$ and $\bm{\Phi}^{\dagger}$ in such a way that
$\bm{\Psi}=U\bm{\Phi},$ (45)
we can write $\hat{H}$ in diagonal form
$\hat{H}=\bm{\Psi}^{\dagger}H\bm{\Psi}=\bm{\Phi}^{\dagger}U^{\dagger}HU\bm{\Phi}=\bm{\Phi}^{\dagger}E_{\text{diag}}\bm{\Phi}.$ (46)
Similarly to $\bm{\Psi}$, we can define new fermion operators $\hat{\gamma}$ such that
$\bm{\Phi}=\begin{pmatrix}\hat{\gamma}\\ \hat{\gamma}^{\dagger}\end{pmatrix}=U^{\dagger}\bm{\Psi}=\begin{pmatrix}U^{\dagger}&V^{\dagger}\\ V^{T}&U^{T}\end{pmatrix}\begin{pmatrix}\hat{\mathbf{c}}\\ \hat{\mathbf{c}}^{\dagger}\end{pmatrix}$. (47)
More explicitly, we can write:
$\begin{cases}\hat{\gamma}_{\mu}=\sum_{j=1}^{L}\left(U^{*}_{j\mu}\hat{c}_{j}+V^{*}_{j\mu}\hat{c}^{\dagger}_{j}\right),\\ \hat{\gamma}^{\dagger}_{\mu}=\sum_{j=1}^{L}\left(V_{j\mu}\hat{c}_{j}+U_{j\mu}\hat{c}^{\dagger}_{j}\right),\end{cases}$ (48)
which can be easily inverted, remembering that $\bm{\Psi}=U\bm{\Phi}$, to express the $\hat{c}_{j}$ operators in terms of the $\hat{\gamma}_{\mu}$:
$\begin{cases}\hat{c}_{j}=\sum_{\mu}\left(U_{j\mu}\hat{\gamma}_{\mu}+V^{*}_{j\mu}\hat{\gamma}^{\dagger}_{\mu}\right),\\ \hat{c}^{\dagger}_{j}=\sum_{\mu}\left(V_{j\mu}\hat{\gamma}_{\mu}+U^{*}_{j\mu}\hat{\gamma}^{\dagger}_{\mu}\right).\end{cases}$ (49)
Finally, $\hat{H}$ in terms of the $\hat{\gamma}$ operators reads, assuming we have taken $\epsilon_{\mu}>0$:
$\hat{H}=\sum_{\mu=1}^{L}\left(\epsilon_{\mu}\hat{\gamma}^{\dagger}_{\mu}\hat{\gamma}_{\mu}-\epsilon_{\mu}\hat{\gamma}_{\mu}\hat{\gamma}^{\dagger}_{\mu}\right)=\sum_{\mu=1}^{L}2\epsilon_{\mu}\left(\hat{\gamma}^{\dagger}_{\mu}\hat{\gamma}_{\mu}-\frac{1}{2}\right),$ (50)
Appendix C Adiabatic Transformation